OpenStudy (anonymous):

how to simplify 1/2(g-16)

OpenStudy (compassionate):

Use the distribution property. Multiply the outside by everything inside the brackets: That is... 1/2 is the same as: 2(g - 16) Since it's multiplication perform the inverse operation: Divide by the terms on the inside: That is g - 16 ------ 2

OpenStudy (compassionate):

More specifically, you multiple the terms by 1/2, giving you. That is you distribute: 1/2 * g = 1g/2 1/2 * (1)-16 = -16 /2 The 1's cancel, leaving you with g - 16 ------ 2

OpenStudy (compassionate):

Any further questions?